OpenStudy (timaashorty):

Given f(x)=x^2-8 and g(x)= (x+1)/x, find (g°f) (1)

OpenStudy (solomonzelman):

to find (fog) multiply what the f(x) AND g(x) together.

OpenStudy (solomonzelman):

(x²-8) × [ (x+1)/x ]

OpenStudy (solomonzelman):

(x²-8)(x+1) --------- x x³ + x² - 8x - 8 -------------- x

OpenStudy (solomonzelman):

to finis the problem, plug in 1 for x. x³ + x² - 8x - 8 (into -------------- ) x

OpenStudy (solomonzelman):

(I can use latex this or next time if you want) Let me know if you \(\Large\color{ blue }{\Large {\bbox[5pt, white ,border:2px solid white ]{ need~~~help? \rm{\Huge✍} }}}\)

OpenStudy (timaashorty):

I got eight

OpenStudy (xapproachesinfinity):

Is it ftimes g or fog?

OpenStudy (timaashorty):

I believe it's fog. because it had that degree symbol not *

OpenStudy (timaashorty):

Solomon do I plug in 1 for all x's??

OpenStudy (timaashorty):

@SolomonZelman

OpenStudy (timaashorty):

these are my answers choices a. 8/7 b. 8 c. 6/7 d. -8 and by solving I got 8. Am I correct?

OpenStudy (solomonzelman):

by plugging in 1, i et 1 + 1 - 8 - 8 ---------- 1 = -14/1 But that is not an Upton -:(

OpenStudy (solomonzelman):

I did (f×g)... oops

OpenStudy (timaashorty):

ooo no wonder why i was confused.

OpenStudy (solomonzelman):

f(x)=x²-8 g(x)= (x+1)/x, find (g°f) (1) (g° f)(x)= ( [x²-8] +1)/ [ x²-8] (g° f)(1)= ( [1²-8] +1)/ [ 1²-8] (g° f)(1)= ( -7 +1)/ -7 (g° f)(1)=-6 / -7 = 6/7 Option C

OpenStudy (solomonzelman):

Sorry for my mistake.

OpenStudy (timaashorty):

It's fine(: Thanks you so much for the help!

OpenStudy (solomonzelman):

Anytime ! (I was offline, sorry for late reply)